# Graph-based Confidence Calibration for Large Language Models

### v1

@@ 4 Experiments
The goal of this section is to compare our proposed framework with baseline methods in terms of confidence calibration. All experiments are conducted on NVIDIA A100 GPUs with 80GB of memory. The supplementary materials and Appendix provide the code for our model, more experiment details in Appendix A.1, and prompting strategy and Appendix A.6.
@@ 4.1 Dataset and Experiment setup
Dataset: We conduct experiments on two public benchmark datasets: (1) CoQA (Reddy et al., 2019), an open-book conversational question answering (QA) task; (2) TriviaQA (Joshi et al., 2017), a commonsense QA task. We use the first 4k as training data and 1582 for validation and testing, both of equal size. We provide more details about training in the Appendix A.1.
Baselines: We compare our methods with the following baselines.
@@
Length-normalized sequence likelihoods (Seq. likelihood) (Malinin & Gales, 2021; Kuhn et al., 2022) is a standard measure for confidence. This method calculates the likelihood of each sequence and normalizes it by the length of the sequence to provide a fair comparison between different lengths of sequences. Platt scaling (Platt, 1999), a variant of the sequence likelihood baseline, applies Platt scaling to the raw likelihoods. GraphSpectral (Chen et al., 2024) uses the graph theory to estimate the confidence. Then we also include post-hoc uncertainty calibration,GraphSpectral+Iso and GraphSpectral+Platt into the baseline methods. Self-check GPT(Manakul et al., 2023) checked the consistency between responses querying the LLM. Verbalized Uncertainty (Lin et al., 2022; Tian et al., 2023; Xiong et al., 2024) generates verbal statements about the model’s confidence in its predictions. Verbalized Qual maps the confidence percent (Verbalized %) into numerical values. APRICOT (Ulmer et al., 2024), a supervised method, fine-tunes the Deberta language model to predict confidence scores for LLM outputs. Furthermore, we also include the baseline of applying two post-hoc uncertainty calibration methods, APRICOT+Iso and APRICOT+Platt, to adjust the confidence scores obtained by Apricot. We performed all the baseline experiments utilizing the open-source codebase.
@@
Graph construction: For each question, we prompt the LLM to generate 30 answers. Each generated answer is then processed using the SentenceBert model Reimers & Gurevych (2019) to obtain the answer’s high-dimensional embeddings. To quantify the semantic similarity between the answers, we compute the cosine similarity between every pair of answer embeddings. These similarity scores are then utilized as edge weights in our similarity graph, where each node represents an individual answer, and the edges signify the degree of semantic relation between them.
@@
Model hyper-parameters: To ensure our model can capture complex and abstract features at each layer, our model comprises three Graph Convolutional Network (GCN) layers, with embedding dimensions of 256, 512, and 1024 for the first, second, and third layers, respectively. The initial learning rate was set to $10^{-4}$. If the validation loss did not show improvement over ten consecutive epochs, the learning rate was reduced by a factor of 0.9. The optimization was performed using the Adam optimizer, configured with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.98$. The batch size was 16.
@@
LLMs: We assess our confidence calibration method on two LLMs with excellent performance: Llama3-8B (Llama3)(Meta, 2024), and Vicuna-7b-v1.5 (Vicuna) (Zheng et al., 2024).
Labeling the data: To obtain the correctness label for CoQA and TriviaQA datasets, we followed previous work (Kuhn et al., 2022) and used the Rougel-L metric for labeling.
@@
Evaluation metrics: The evaluation metrics include Expectation Calibration Error (ECE), Brier Score, and AUROC. Specifically, (1) ECE quantifies the consistency between the prediction error and the uncertainty of the prediction. An ideal calibration curve should exhibit a lower ECE. It measures the consistency between the prediction error and the confidence of the prediction. Specifically, the confidence interval is grouped into fixed bins, and the average of the difference between the confidence and error in each bin is compared. Formally, ECE is calculated as
$ECE=\sum_{b=1}^{B}\frac{n_{b}}{N}|acc(b)-conf(b)|$ (8)
where $n_{b}$ is the number of predictions in bin $b$, $N$ is the total number of data points and $acc(b)$ and $conf(b)$ are the accuracy and confidence of bin b, respectively. (2) Brier Score (Brier, 1950), which is the mean squared difference between predicted probabilities and the actual binary results. Lower Brier Scores indicate better performance. (3) AUROC to indicate the models’ discriminatory ability.
@@
Table 1: Comparison of confidence calibration performance on TriviaQA and CoQA dataset for Llama3
Method TriviaQA CoQA
@@
GraphSpectral (GS) 0.22 0.80 0.056 0.19 0.76 0.091
GS + Iso 0.16 0.80 0.072 0.16 0.76 0.054
GS + Platt 0.16 0.79 0.048 0.16 0.74 0.041
Self-check GPT 0.33 0.65 0.187 0.21 0.63 0.178
Seq. likelihood 0.54 0.53 0.22 0.38 0.47 0.17
Platt 0.28 0.59 0.05 0.26 0.47 0.09
Verbalized Qual 0.32 0.62 0.14 0.30 0.68 0.16
Verbalized % 0.25 0.67 0.033 0.42 0.66 0.21
APRICOT 0.14 0.76 0.074 0.17 0.78 0.132
APRICOT + Iso 0.18 0.76 0.073 0.17 0.78 0.097
APRICOT + Platt 0.17 0.76 0.039 0.17 0.78 0.069
Ours 0.14 0.84 0.025 0.12 0.77 0.016
Ours (Multi prompts) 0.14 0.84 0.025 0.12 0.78 0.015
4.2 Experiment Results
For the Llama3 model, the confidence calibration performance on TriviaQA and CoQA are shown in Table 1. For the TriviaQA dataset, it can be observed that the likelihood-based method performs poorly on the calibration error (ECE and Brier Score) and AUROC due to unreliable model prediction probability (Zhang et al., 2024). Platt scaling improves the ECE post-calibration and enhances the model’s discriminative ability, resulting in higher AUROC results. However, this method cannot capture the semantic equivalence among answers, leading to sub-optimal performance. The Verbalized and Verbalized Qual prompts LLM to output confidence for their answers, improving AUROC by $3-5\%$ compared with the likelihood baseline. However, it faces the overconfidence issue; thus, the calibration errors are still high. The graphSpectral method can produce good confidence estimations, but its calibration performance is poor. Even with the addition of techniques such as Isotonic Calibration or Platt Scaling, this issue can only be partially mitigated. The auxiliary DeBERTa method combines the LLM outputs, Chain-of-Thoughts (CoT) outputs, and verbalized confidence to fine-tune the DeBERTa model for predicting confidence. Our method captures the prediction confidence based on the graph structure of LLM’s responses in semantic space and achieves better ECE and smECE results. The ECE is reduced from $0.07$ to $0.02$ and improves the AUROC from $0.76$ to $0.84$ compared with the baseline calibration methods. The experiment results on CoQA for the Llama3 model are shown in Table 1. These results show a similar trend, with our model achieving superior performance in confidence calibration compared to the baseline methods.
Table 2: Comparison of confidence calibration performance on TriviaQA and CoQA dataset for Vicuna
Method TriviaQA CoQA
@@ 4.2 Experiment Results
GraphSpectral (GS) 0.19 0.79 0.112 0.27 0.69 0.202
GS + Iso 0.19 0.79 0.059 0.24 069 0.037
GS + Platt 0.17 0.79 0.067 0.22 0.69 0.055
Self-check GPT 0.35 0.63 0.18 0.22 0.64 0.192
Verbalized Qual 0.38 0.62 0.02 0.45 0.48 0.00
Verbalized % 0.39 0.52 0.38 0.49 0.53 0.32
Seq. likelihood 0.47 0.58 0.42 0.30 0.68 0.16
Platt 0.34 0.58 0.25 0.30 0.68 0.16
APRICOT 0.18 0.78 0.068 0.19 0.74 0.073
APRICOT + Iso 0.17 0.78 0.049 0.19 0.74 0.06
APRICOT + Platt 0.17 0.78 0.052 0.19 0.74 0.049
Ours 0.17 0.81 0.020 0.18 0.75 0.02
Ours (Multi prompts) 0.16 0.81 0.018 0.16 0.75 0.016
Furthermore, we also compare the confidence calibration performance for the Vicuna model on the TriviaQA and CoQA datasets. The results are summarized in Table 2. Our model consistently improves the calibration error compared to the baseline methods.
To understand the improvements of our model, we present the reliability diagrams for the baseline methods applied to the Vicuna model on TriviaQA. The reliability diagram is created by discretizing the confidence value into $10$ bins and then computing the average accuracy for samples in each bin. The ideal calibration curve should align with the diagonal line, indicating that the confidence value can match the probability of correctness. The reliability diagram is shown in Fig. 2. (We also show other reliability diagrams for the different methods for Llamas on TriviaQA and CoQA in the Appendix  A.5). The figure presents the reliability diagrams for different methods, each utilizing 10 bins. In these diagrams, both the color intensity and the percentage numbers within each bar represent the proportion of total responses that fall into each respective bin. Specifically, larger proportions are depicted with colors closer to purple, while the height of each bar indicates the ratio of correct predictions within that bin. An ideal reliability diagram should exhibit a wide distribution of responses across multiple bins, demonstrating the model’s strong ability to differentiate between varying confidence levels in its predictions. Additionally, the heights of the bars should align closely with the diagonal line, which represents perfect calibration—where the predicted confidence matches the empirical accuracy. It can be observed that the likelihood-based confidence methods exhibit significant overconfidence, with curves below the diagonal, indicating many samples have high confidence but low accuracy. This results in poor ECE performance. Although the Platting scaling calibration method enhances the ECE performance, it still has poor AUROC. The Auxiliary DeBERTa (APRICOT) method, which integrates LLM outputs, Chain-of-Thought (CoT) outputs, and verbalized confidence to train an auxiliary DeBERTa model, enhances the AUROC. However, it still experiences some overconfidence issues, potentially caused by the inherent overconfidence in the input verbalized confidence scores. The baseline method’s reliability diagrams revealed that this method frequently assigned high confidence scores to incorrect predictions, deviating markedly from the ideal calibration represented by the diagonal line. For example, the verbalized method’s predictions in the highest confidence bins (80-90%) were significantly below the corresponding empirical accuracy, indicating a tendency to overestimate the certainty of its outputs. In contrast, our framework achieves a broad spread of responses across the bins, showing good differentiation capabilities; at the same time, the bar heights closely follow the diagonal line, indicating better calibration.
@@
Figure 2: Reliability diagrams for different methods using 10 bins each for Vicuna on TriviaQA. The color, as well as the percentage number within each bar, indicates the proportion of total responses contained in each bin. Larger values are represented by colors closer to purple, and the height indicates the ratio of correct ones. We prefer a wide spread of responses in different bins (strong ability to differentiate responses) and bin heights along the diagonal line (accurate calibration). In comparison to all competing methods, our model consistently outperforms others by exhibiting a more extensive spread across the bins and maintaining calibration accuracy along the diagonal.
4.3 Out of Domain evaluation

### v2

@@ 4 Experiments
We evaluate the calibration performance of our proposed framework by comparing it against baseline methods. We also test their performances in the OOD setting, where their calibration parameters are decided on one dataset, but they are tested on a different dataset. We further analyze the quality of automatic labels and the sensitivities of our method under different hypter-parameter settings.
@@ 4.1 Dataset and experiment setup
Dataset: We conduct experiments on four datasets: (1) CoQA (Reddy et al., 2019), an open-book conversational question answering task; (2) TriviaQA (Joshi et al., 2017), a commonsense QA task. (3) TruthfulQA (Lin et al., 2022a), a comparably more challenging dataset for factual QA tasks. and (4) HotpotQA, a question answering dataset that requires models to find and combine information from multiple passages to answer complex questions. We repeat the experiments 10 times, each with a different train/validation split and test the performance on the test set.
@@
Baselines: We compare our methods with the following baselines. Length-normalized sequence likelihoods (Seq. likelihood(Malinin & Gales, 2021; Kuhn et al., 2022) is a standard measure for confidence. This method calculates the likelihood of each sequence and normalizes it by the length of the sequence to provide a fair comparison between different lengths of sequences. Platt scaling (Platt, 1999), a variant of the sequence likelihood baseline, applies Platt scaling to the raw likelihoods. GraphSpectral (Lin et al., 2024) uses the graph theory to estimate the confidence. Then we also include post-hoc uncertainty calibration, GraphSpectral+Iso and GraphSpectral+Platt into the baseline methods. Self-check GPT(Manakul et al., 2023) checked the consistency between responses querying the LLM. Verbalized Uncertainty (Lin et al., 2022b; Tian et al., 2023; Xiong et al., 2024) generates verbal statements about the model’s confidence in its predictions. Verbalized Qual maps the confidence percent (Verbalized %) into numerical values. APRICOT (Ulmer et al., 2024), a supervised method, fine-tunes the Deberta language model to predict confidence scores for LLM outputs. Furthermore, we also include the baseline of applying two post-hoc uncertainty calibration methods, APRICOT+Iso and APRICOT+Platt, to adjust the confidence scores obtained by Apricot. We performed all the baseline experiments utilizing the open-source codebase and used the default hyperparameters.
@@
Graph construction: For each question, we generate 30 candidate answers through LLM prompting. Each generated response is then processed using the Sentence-BERT (SBERT) model (Reimers & Gurevych, 2019) to obtain the answer’s high-dimensional embeddings. To quantify the semantic alignment between all responses, we compute the cosine similarity between every pair of answer embeddings. These scores are then utilized as edge weights in our similarity graph, where each node represents an individual answer, and the edges weights signify the degree of semantic relation between them.
@@
Model hyper-parameters: To ensure our model can capture complex and abstract features at each layer, our model comprises three GNN layers, with embedding dimensions of 256, 512, and 1024 for the first, second, and third layers, respectively. The initial learning rate was set to $10^{-4}$. If the validation loss did not show improvement over ten consecutive epochs, the learning rate was reduced by a factor of 0.9. The optimization was performed using the Adam optimizer, configured with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.98$.
@@
LLMs: We assess our confidence calibration method on two LLMs with good performance: Llama3-8B (Llama3)(Meta, 2024), and Vicuna-7b-v1.5 (Vicuna) (Zheng et al., 2024).
Labeling the data: To obtain the correctness label for CoQA and TriviaQA datasets, we followed previous work (Kuhn et al., 2022) and used the ROUGE-L metric for labeling. For the TruthfulQA dataset, given its focus on factual correctness and longer answers, we employed GPT4 (Potsaweel, 2024; Liu et al., 2023; Badshah & Sajjad, 2024) to generate the labels. In addition, we conducted experiments on a smaller subset of data using human-annotated labels to validate the reliability of our automatic labeling process. Details of this manual annotation study are provided in Section4.4.
@@
Evaluation metrics: The evaluation metrics include Expectation Calibration Error (ECE), Brier Score, and AUROC. Specifically, (1) ECE quantifies the consistency between the prediction error and the uncertainty of the prediction. An ideal calibration curve should exhibit a lower ECE. It measures the consistency between the prediction error and the confidence of the prediction. Specifically, the confidence interval is grouped into fixed bins, and the average of the difference between the confidence and error in each bin is compared. Formally, ECE is calculated as $ECE=\sum_{b=1}^{B}\frac{n_{b}}{N}|acc(b)-conf(b)|,$ where $n_{b}$ is the number of predictions in bin $b$, $N$ is the total number of data points and $acc(b)$ and $conf(b)$ are the accuracy and confidence of bin b, respectively. (2) Brier Score (Brier, 1950), which is the mean squared difference between predicted probabilities and the actual binary results. Lower Brier Scores indicate better performance. (3) AUROC to indicate the models’ discriminatory ability.
Further experimental configurations and prompting strategy are provided in the Appendix. The experiments are conducted on NVIDIA A100 GPUs with 80GB of memory.
4.2 Experiment results
For the Llama3 model, the confidence calibration performance on TriviaQA is shown in Table 1. For the TriviaQA dataset, it can be observed that the likelihood-based method performs poorly on the calibration error (ECE and Brier Score) and AUROC due to unreliable model prediction probability (Zhang et al., 2024). Platt scaling improves the ECE post-calibration and enhances the model’s discriminative ability, resulting in higher AUROC results. However, this method cannot capture the semantic equivalence among answers, leading to sub-optimal performance. The Verbalized and Verbalized Qual prompts LLM to output confidence for their answers, improving AUROC by $3-5\%$ compared with the likelihood baseline. However, it faces the overconfidence issue; thus, the calibration errors are still high. The GraphSpectral method can produce good confidence estimations, but its calibration performance is poor. Even with the addition of techniques such as Isotonic Calibration or Platt Scaling, this issue can only be partially mitigated. The auxiliary DeBERTa method combines the LLM outputs, Chain-of-Thoughts (CoT) outputs, and verbalized confidence to fine-tune the DeBERTa model for predicting confidence. Our method captures the prediction confidence based on the graph structure of LLM’s responses in semantic space and achieves better ECE results. The ECE is reduced from $0.07$ to $0.03$ and improves the AUROC from $0.72$ to $0.86$ compared with the baseline calibration methods. The experiment results on TruthfulQA, HotpotQA and CoQA for the Llama3 model are shown in Table 1. These results show a similar trend, with our model achieving superior performance in confidence calibration compared to the baseline methods.
@@ 4.2 Experiment results
Table 1: Comparison of confidence calibration performance on TriviaQA, CoQA, TruthfulQA and HotpotQA dataset for Llama3
Method TriviaQA CoQA TruthfulQA HotpotQA
@@
GraphSpectral (GS) .223 ± .002 .842 ± .002 .0762 ± .007 .193 ± .001 .762 ± .008 .110 ± .019 .332 ± .002 .667 ± .012 .239 ± .019 .172 ± .017 .783 ± .006 .097 ± .014
GS + Iso .167 ± .011 .842 ± .002 .058 ± .002 .162 ± .008 .762 ± .008 .054 ± .002 .191 ± .015 .667 ± .012 .088± .007 .163 ± .012 .783 ± .006 .087 ± .021
GS + Platt .165 ± .012 .842 ± .002 .049 ± .002 .161 ± .009 .762 ± .008 .042 ± .001 .221 ± .013 .667 ± .012 .151± .008 .160 ± .014 .783 ± .006 .177 ± .012
Self-checkGPT .332 ± .031 .652 ± .020 .187 ± .002 .209 ± .020 .633 ± .027 .178 ± .010 .362 ± .028 .566 ± .028 .353± .030 .283 ± .014 .673 ± .022 .122 ± .030
Seq. likelihood .536± .015 .591 ± .002 .220 ± .002 .382 ± .012 .571 ± .028 .173 ± .009 .465 ± .008 .582 ± .025 .052 ± .009 .463 ± .018 .651 ± .002 .105± .012
Platt .276 ± .006 .591 ± .002 .052 ± .002 .258 ± .000 .571 ± .0280 .090 ± .009 .271 ± .007 .582 ± .025 .053 ± .008 .220 ± .0120 .651 ± .002 .142± .008
Verbalized Qual .322 ± .034 .618 ± .002 .142 ± .002 .302 ± .021 .681 ± .022 .160 ± .007 .320 ± .037 .622 ± .016 .140 ± .008 .358 ± .012 .652 ± .006 .150 ± .029
Verbalized % .253 ± .021 .663 ± .008 .033 ± .002 .423 ± .0120 .662 ± .027 .216 ± .002 .540 ± .035 .573 ± .029 .331 ± .008 .319 ± .014 .672 ± .002 .220± .023
APRICOT .145 ± .002 .723 ± .003 .074 ± .005 .173 ± .006 .751 ± .022 .132 ± .006 .201 ± .003 .657 ± .034 .062 ± .011 .171 ± .014 .823 ± .011 .081± .009
APRICOT+Iso .182 ± .012 .723 ± .003 .073 ± .004 .171 ± .009 .751 ± .022 .097 ± .003 .200 ± .003 .657 ± .034 .059 ± .011 .180 ± .012 .823 ± .011 .073± .002
APRICOT+Platt .173 ± .018 .723 ± .003 .042 ± .004 .169 ± .012 .751 ± .022 .069 ± .008 .230 ± .003 .657 ± .034 .056 ± .011 .171 ± .018 .823 ± .011 .071± .010
Ours .136 ± .000 .864 ± .002 .035 ± .004 .124 ± .000 .768 ± .009 .013 ± .003 .151 ± .003 .732 ± .012 .037±.013 .142 ± .000 .815 ± .002 .023 ± .004
Ours(Multi prompts) .141 ± .002 .853 ± .002 .036 ± .008 .118 ± .000 .776 ± .012 .015 ± .007 .173 ± .003 .736 ± .007 .039 ± .013 .142 ± .000 .821 ± .002 .021 ± .006
Table 2: Comparison of confidence calibration performance on TriviaQA, CoQA, TruthfulQA and HotpotQA dataset for Vicuna
Method TriviaQA CoQA TruthfulQA HotpotQA
@@
GraphSpectral (GS) .196±.000 .792±.006 .112±.014 .275±.004 .696±.004 .202±.011 .286±.006 .647±.009 .226±.013 .162±.076 .673±.008 .202±.009
GS + Iso .196±.000 .792±.006 .059±.008 .245±.002 .696±.004 .037±.014 .297±.008 .647±.009 .092±.004 .165±.058 .673±.008 .085±.028
GS + Platt .172±.000 .792±.006 .067±.009 .228±.002 .696±.004 .055±.028 .307±.007 .647±.009 .183±.003 .160±.049 .673±.008 .073±.049
Self-checkGPT .355±.001 .640±.003 .183±.014 .221±.004 .648±.009 .192±.010 .281±.012 .552±.008 .308±.017 .295±.187 .652±.187 .370±.187
Seq. likelihood .485±.002 .581±.002 .420±.029 .302±.012 .688±.002 .169±.090 .325±.022 .587±.010 .205±.012 .493±.220 .630±.050 .223±.220
Platt .342±.002 .581±.002 .255±.016 .308±.015 .688±.002 .165±.008 .288±.014 .587±.010 .181±.005 .232±.050 .630±.050 .259±.050
Verbalized Qual .393±.002 .631±.007 .029±.014 .455±.022 .495±.004 .009±.001 .471±.034 .482±.060 .018±.005 .220±.140 .652±.140 .142±.140
Verbalized % .402 ±.001 .523±.005 .383±.012 .492±.025 .539±.003 .324±.029 .580±.022 .566±.009 .387±.017 .342±.033 .683±.033 .033±.033
APRICOT .196±.000 .783±.006 .068±.007 .193±.004 .742±.006 .073±.009 .197±.007 .769±.002 .118±.005 .152±.001 .782±.022 .074±.074
APRICOT+Iso .187±.000 .783±.006 .049±.004 .193±.005 .742±.006 .064±.007 .197±.007 .769±.002 .092±.008 .142±.001 .782±.022 .073±.073
APRICOT+Platt .186±.000 .783±.006 .052±.004 .193±.005 .742±.002 .049±.004 .204±.006 .769±.002 .085±.005 .150±.001 .782±.022 .042±.042
Ours .169±.000 .816±.002 .028±.004 .184±.001 .754±.004 .032±.004 .202±.003 .774±.001 .059±.006 .132±.000 .791±.002 .022±.002
Ours(Multi prompts) .165±.000 .815±.006 .025±.003 .168±.001 .763±.004 .030±.006 .202±.004 .764±.001 .063±.004 .131±.000 .790±.003 .025±.009
Furthermore, we also compare the confidence calibration performance for the Vicuna model on the TriviaQA, CoQA, TruthfulQA and HotpotQA datasets. The results are summarized in Table 2. Our model consistently improves the calibration error compared to the baseline methods. Both GraphSpectral and our method have a similar assumption that the consistency level between responses indicates the confidence levels of these responses. However, GraphSpectral uses simple graph statistics to measure the confidence level of responses and could not capture complex relationships between response patterns and confidence levels (e.g. patterns beyond clustering structures). As a comparison, by framing the problem as a learning problem, our method has better opportunities to discover such relationships and provides a better calibration performance. Self-Check GPT uses its own evaluation on whether the context supports the answers and heavily relies on the LLM model’s capability to do self-reflection, which can also be hallucinated. Thus the generated confidence scores are not calibrated well with empirical accuracy.
We present the reliability diagrams for all methods on TriviaQA to better understand the model improvement. The reliability diagram is created by binning responses into $10$ bins according to their confidence values. Then the frequency of correctness is also computed for each bin. With an ideal calibration, the confidence value in each bin should match the frequency of correctness. At the same time, a good calibration model should differentiate responses according to their confidence levels and give a wide distribution of confidence values in these bins.
The reliability diagram is shown in Fig. 2. (We also show other reliability diagrams for the different methods for Llamas on TriviaQA and CoQA in the Appendix  E). The figure presents the reliability diagrams for different methods. In these diagrams, both the color intensity and the percentage numbers within each bar represent the proportion of total responses that fall into each respective bin. Specifically, larger proportions are depicted with colors closer to purple, while the height of each bar indicates the ratio of correct predictions within that bin. Our framework achieves a broad spread of responses across the bins, showing good differentiation capabilities; at the same time, the bar heights closely follow the diagonal line, indicating a good calibration performance.
@@
Figure 2: Reliability diagrams for different methods using 10 bins each for Vicuna on TriviaQA. The color, as well as the percentage number within each bar, indicates the proportion of total responses contained in each bin. Larger values are represented by colors closer to purple, and the height indicates the ratio of correct ones. We prefer a wide spread of responses in different bins (strong ability to differentiate responses) and bin heights along the diagonal line (accurate calibration). Our model outperforms others with a broader bin spread and better alignment with the diagonal for calibration accuracy.
However, baseline methods cannot reach the same performance. The likelihood-based confidence methods exhibit significant overconfidence, indicating many responses rated with high confidence are actually wrong. The Auxiliary DeBERTa (APRICOT) method, which integrates LLM outputs, Chain-of-Thought (CoT) outputs, and verbalized confidence to train an auxiliary DeBERTa model, enhances the AUROC. However, it still experiences some overconfidence issues, potentially caused by the inherent overconfidence in the input verbalized confidence scores. Furthermore, the baseline methods’ reliability diagrams revealed that this method frequently assigned high confidence scores to incorrect predictions, deviating markedly from the ideal calibration represented by the diagonal line. For example, the verbalized method’s predictions in the highest confidence bins (80-90%) were significantly below the corresponding empirical accuracy, indicating a tendency to overestimate the certainty of its outputs.
4.3 Confidence calibration in the OOD setting